A resistor dissipates 100W of power when it has a 10A current through it. What is the current through the same resistor when connected to a 20V emf

please show work
Answer #1

I=10A

P=I^2R

100 = (10)^2R

R=1ohm

emf =20V

I=emf/R = 20/1

I=20A
Answer #2

resistance value=power/i^2=100/10^2= 1 ohm

current when 20 v connected= v/R=20/1 =20 A
Answer #3

P = VI
V = P/V = 100/10 = 10 V

R = V/I =10/10 = 1 ohm

I = V/R = 20/1 = 20 A